What is the new freezing point of a glucose (C6H12O6) solution if 150 g of glucose is dissolved in 500 g of water?


1
Expert's answer
2021-03-29T06:02:03-0400

Molar Mass of C6H12O6 = 180.169 g/mol

Molar Mass of H2O = 18.01528 g/mol

N of moles C6H12O6 = 150/180.16

= 0.833 M

Mass Kg of water = 0.5 Kg

Molality = 0.8/0.5 = 1.6 mol/Kg

H2O has 2 ion ; H+ and (OH)-

i = 2

Freezing point of water = 0°C

Kb (°C/m) = 0.52

= 0.52×2×1.6

= 1.664+ 0

= 1.664

= 1.66°C